Pranks, Inc. Pranks, Inc. is a manufacturer of jok
Answer #1

Part 1)

The number of common stock outstanding, preferred stock outstanding and preferred dividend as a percent of par is calculated as follows:

Number of Common Stock Outstanding = Total Common Stock Dividend Paid in 2016/Dividend Paid Per Share = 150,000/.30 = 500,000 shares

______

Number of Preferred Stock Outstanding = Total Preferred Dividend Paid in 2016/Dividend Paid Per Share = 60,000/.30 = 200,000 shares

______

Preferred Dividend as a Percent of Par = (Preferred Dividend Per Share)/Par Value*100 = .30/30*100 = 1%

______

Part 2)

The value of paid-in capital before stock dividend, total retained earnings before stock dividend and total stockholder's equity before stock dividend is given below:

Total Paid-in Capital (Common Stock) [500,000*20 + 7,000,000] 17,000,000
Total Paid-in Capital (Preferred Stock) [200,000*30] 6,000,000
Total Paid-in Capital Before Stock Dividend $23,000,000
Retained Earnings before Stock Dividend $33,500,000
Total Stockholder's Equity before Stock Dividend [23,000,000 + 33,500,000] $56,500,000

_____

The value of paid-in capital after stock dividend, total retained earnings after stock dividend and total stockholder's equity after stock dividend is given below:

Total Paid-in Capital (Common Stock) [500,000*20 + 7,000,000 + 500,000*2%*20] 17,200,000
Total Paid-in Capital (Preferred Stock) [200,000*30] 6,000,000
Total Paid-in Capital after Stock Dividend $23,200,000
Retained Earnings after Stock Dividend [33,500,000 - 500,000*2%*24] $33,260,000
Total Stockholder's Equity before Stock Dividend [23,000,000 + 33,500,000 + (240,000 - 200,000)] $56,500,000

______

Part 3)

1)

Yes, because the number of shares issued (750,000) is greater than the number of shares outstanding (500,000). (which is Option B)

______

2)

The annual preferred dividend is $60,000 (200,000*.30). The company paid a dividend of $40,000 in Year 2011. The balance amount of $20,000 (200,000*.30 - 40,000) in arrears for 2011 is paid as follows:

Year 2012 = 72,000 - 60,000 = $12,000

Year 2013 = 68,000 - 60,000 = $8,000